# Quasi-Monte Carlo with one categorical variable

### v2

@@ 2.5 Discrepancy within each stratum
For $S_{\ell}$ defined at (4), sample $i\in S_{\ell}$ holds if and only if $B_{\ell-1}\leqslant z_{i1}<B_{\ell}$ or equivalently
$\sum_{1\leqslant r<\ell}2^{-\kappa_{r}}\leqslant z_{i1}<\sum_{1\leqslant r\leqslant\ell}2^{-\kappa_{r}}$
which we can write as
Now consider a dyadic box in $[0,1]^{s}$ of the form
$\prod_{j=1}^{s}\Bigl{[}\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr{)}.$
@@
Its absolute value is at most
$\Bigl{|}\frac{n\beta_{\ell}}{n}-1\Bigr{|}\mathrm{Vol}([\bm{0},\bm{c}))+\frac{2n}{n_{\ell}}O(n^{-1}\log(n)^{s})$ $=O(n_{\ell}^{-1}\log(n_{\ell}/\beta_{\ell})^{s})$
$=O(n_{\ell}^{-1}\log(n_{\ell})^{s})$
@@ 3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the convergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases. It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer. Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$. Those often appear to have $21<\rho<$ 2.
@@
For $\rho>2$, we have the reverse situation. Changing $n_{\ell}$ by even ±$\pm 1$ can worsen the convergence rate because changing $\hat{\mu}$ by $O(1/n)$ is a bigger change than the RMSE of $O(n^{-\rho/2})$. A convergence rate with $\rho>2$ for scrambled Sobol’ points is only available for sample sizes that are powers of two. Section 4 shows how to obtain $n_{\ell}$ equal to powers of two.
@@ 3.1 Optimal sample sizes
Under the model (9),
${\mathrm{Var}}(\hat{\mu})={\mathrm{Var}}\biggl{(}\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}\biggr{)}\leqslant\biggl{(}\sum_{\ell=1}^{L}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}\biggr{)}^{2}$ (10)
where the upper bound is from the very conservative assumption that the estimates within all $L$ strata have unit correlations. With no prior information on the relative sizes of $\tau_{\ell}$ we adopt
$C_{1}(\rho)=C_{1}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}$ (11)
as a criterion to minimize subject to $\sum_{\ell=1}^{L}n_{\ell}=n$. We form the Lagrangian
$G_{1}=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}+\lambda\sum_{\ell=1}^{L}n_{\ell}$
and setting $\partial G_{1}/\partial n_{\ell}$ to zero yields a solution with
$n_{\ell}\propto\alpha_{\ell}^{2/(\rho+2)}.$ (12)
We then suppose that the user chooses integers $n_{\ell}$ summing to $n=2^{m}$ that are nearly proportional to $\alpha_{\ell}^{2/(\rho+2)}$.
RQMC points have a space filling property. By construction, the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$. We might then anticipate primarily non-positive correlations among the $\hat{\mu}_{\ell}$. This motivates a different criterion
$C_{0}(\rho)=C_{0}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}^{2}n_{\ell}^{-\rho}$ (13)
based on zero correlations among the $\hat{\mu}_{\ell}$. Forming the Lagrangian for this criterion and setting to zero the partial derivatives with respect to $n_{\ell}$ leads to
@@
In both criteria (12) and (14), higher convergence rates $\rho$ tend to equalize the $n_{\ell}$. We can think of $n_{\ell}=n/L$ as arising in the limit as $\rho\to\infty$. We will also see equal $n_{\ell}$ in a minimax argument in Section 5.
We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzes 00 and $1$, and three different hypotheses on $\rho$. They lead to the four different rates given in Table 1. The case for ansatz 00 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation. The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$. This is also true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) do not distinguish between different values of $\rho$. After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
When we do have information about the relative sizes of $\tau_{\ell}$, then we can replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}$ in the formulas in Table 1. In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934). When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$, then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}/c_{\ell}$ in Table 1.
@@
Table 1: This table shows ideal rates for $n_{\ell}$ as functions of $\alpha_{\ell}$. Under ansatz 00, estimates from different strata are uncorrelated. Ansatz $1$ uses a bound assuming that those estimates are perfectly correlated. These rates apply when the variance within stratum $\ell$ decreases proportionally to $n_{\ell}^{-\rho}$.
If $\rho$ varies from stratum to stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$. Strata with higher $\rho_{\ell}$ get larger $n_{\ell}$ and $\beta_{\ell}=n_{\ell}/n$ does not depend on $n$.
@@ 3.2 Suboptimality
In ansatz 0, we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
@@
The factor $n^{-\rho}$ has canceled from the numerator and denominator and so we get the same rate of convergence with an incorrect $\gamma$ as we would with the optimal $\gamma=\rho$. Only the lead constant is affected.
For ansatz 1, assuming a rate $n_{\ell}^{-\gamma}$ instead of $n_{\ell}^{-\rho}$ would lead to $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+2)}$. To put the inefficiency in squared error terms, comparable to $I_{0}$ we use $C^{2}_{1}(\gamma)/C^{2}_{1}(\rho)$, getting
$I_{1}(\gamma\!\mid\!\rho)$ $=\left(\frac{\sum_{\ell=1}^{L}\alpha_{\ell}\bigl{(}n\alpha_{\ell}^{2/(\gamma+2)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr{)}^{-\rho/2}}{\sum_{\ell=1}^{L}\alpha_{\ell}\bigl{(}n\alpha_{\ell}^{2/(\rho+2)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr{)}^{-\rho/2}}\right)^{2}$
$=\frac{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{1-\rho/(\gamma+2)}\bigr{)}^{2}\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr{)}^{\rho}}{\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr{)}^{\rho+2}}.$
We think that ansatz 0 makes a more reasonable assumption than ansatz 1. Then a cautious choice is to take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma_{0}+1)}$ where
@@
Next we consider recommendations that can be made generally without specific values of $\alpha_{\ell}$ in mind. The first thing to notice is that both $I_{0}$ and $I_{1}$ contain a numerator sum that could potentially have a negative power of $\alpha_{\ell}$. For general use where some $\alpha_{\ell}$ might be small, we want to avoid that possibility. There is no possibility of an even more negative power in the denominator to compensate. To keep a positive exponent in the numerator of $I_{0}$ we must have $2-2\rho/(\gamma+1)\geqslant 0$. After rearrangement this becomes a constraint that $\gamma\geqslant\rho-1$, whatever the unknown $\rho$ might be. When $\rho\doteq 3$ is a possibility, then we should take $\gamma\geqslant 2$. If instead we use $\gamma=1$ in a $\rho=3$ setting then the first factor in $I_{0}$ is greater than $\alpha_{L}^{-1}$. For $I_{1}$ we want $1-\rho/(\gamma+2)\geqslant 0$ or $\gamma\geqslant\rho-2$. Since we already had $\gamma\geqslant\rho-1$ using ansatz 0, this is not a further constraint. When we know that $\rho\leqslant 2$, then any $\gamma\geqslant 1$ avoids a negative exponent for $\alpha_{\ell}$.
@@ 4 Making each $n_{\ell}$ a power of two
Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to a given total sample size $n=2^{m}$. We assume throughout that $n\geqslant L$ and that we want every $n_{\ell}\geqslant 1$. The advantage of having every $n_{\ell}$ be a power of two is that we then get scrambled net sampling within each of the $L$ strata as shown by Proposition 2. Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen, 2020). The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
@@
The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc., 2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al. (2023)). The latter reference mentions that the number of solutions is asymptotic to $R\theta^{L}$ for $R\doteq 0.14$ and $\theta\doteq 1.794$. This is a somewhat faster growth rate than the Fibonacci numbers have: their counterpart to $\theta$ is $(1+\sqrt{5})/2\doteq 1.618$. For $L=10$, the approximation is close to 48.3.
@@
Under ansatz 0, the ideal proportion of samples in stratum $\ell$ is
@@
which we consider to be the most ‘under-represented’ of the eligible strata. The analogous algorithm under ansatz $1$ is to double the sample size of stratum
@@
The pseudo-code for forward allocation algorithm for ansatz 0 is given in Algorithm 1.
@@ 5 Minimax choice
$\Delta_{N}$ =$=\biggl{\{}\bm{n}=(n_{1},\dots,n_{L})\in\mathbb{N}^{L}\Bigm{|}\sum_{\ell=1}^{L}n_{\ell}=N\biggr{\}},\quad\text{and}\quad$
$\mathcal{T}$ =$=\Bigl{\{}\bm{\tau}=(\tau_{1},\dots,\tau_{L})\in[0,\infty)^{L}\bigm{|}\max_{1\leqslant\ell\leqslant L}\tau_{\ell}>0\Bigr{\}}$
@@
We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used. For ansatzes 00 and $1$ we define suboptimality ratios
@@
The same argument applies to both ansatzes. We prove it for ansatz 1. Fixing $\bm{n}$ and $\tilde{\bm{n}}$ we see that the maximum over $\bm{\tau}$ of $R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})$ arises when $\tau_{\ell}=0$ for all $\ell$ except $\ell^{*}(\bm{n},\tilde{\bm{n}})=\operatorname*{argmax}_{\ell}(\tilde{n}_{\ell}/n_{\ell})^{\rho/2}$ with $\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}$ then taking any value $\lambda>0$.
Now, we can go back to the initial problem and rewrite (18) as:
after noting that the factor $\alpha_{\ell^{*}(\bm{n},\tilde{\bm{n}})}\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}^{1/2}$ cancels from numerator and denominator. To maximize this ratio over $\tilde{\bm{n}}$ one takes $\tilde{\bm{n}}$ to be the vector with value $N-L+1$ in component $\ell^{*}=\operatorname*{argmin}_{\ell}n_{\ell}$ and ones elsewhere. Then the minimax problem becomes $\min_{\bm{n}\in\Delta_{N}}\bigl{(}(N-L+1)/\min_{1\leqslant\ell\leqslant N}n_{\ell}\bigr{)}^{\rho/2}$. The solution is to maximize the minimum value of $n_{\ell}$. We do that by making $\min_{\ell}n_{\ell}=\lfloor N/L\rfloor$. ∎

### v3

@@ 2.5 Discrepancy within each stratum
[…]
@@
[…]
@@ 3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
Values of $\rho$ equal to $1$, 22 and $3$ correspond to upper bounds on the convergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases. It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer. Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$. Those often appear to have $1<\rho<2$.
@@
For $\rho>2$, we have the reverse situation. Changing $n_{\ell}$ by even $\pm 1$ can worsen the convergence rate because changing $\hat{\mu}$ by $O(1/n)$ is a bigger change than the RMSE of $O(n^{-\rho/2})$. A convergence rate with $\rho>2$ for scrambled Sobol’ points is only available for sample sizes that are powers of two. Section 4 shows how to enforce that all $n_{\ell}$ are equal to powers of two.
@@ 3.1 Optimal sample sizes
Under the model (7), ${\mathrm{Var}}(\hat{\mu})={\mathrm{Var}}\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}\bigr)\leqslant\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}\bigr)^{2}$where the upper bound is from the very conservative assumption that the estimates within all $L$ strata have unit correlations. RQMC points have a space filling property: the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$. Then a zero or even negative correlation among stratum means is more plausible. Under a working model of 0 inter-stratum correlation, ${\mathrm{Var}}(\hat{\mu})\approx\sum_{\ell=1}^{L}\alpha_{\ell}^{2}\tau_{\ell}/n_{\ell}^{\rho}$.
Then with no prior information on the relative sizes of $\tau_{\ell}$ we adopt the criterion
$C_{0}(\rho)=C_{0}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}^{2}n_{\ell}^{-\rho}.$ (8)
The conservative criterion assuming unit correlations is $C_{1}(\rho)=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}$.
Forming the Lagrangian for (8) with the constraint $\sum_{\ell}n_{\ell}=n$ and setting to zero the partial derivatives with respect to $n_{\ell}$ leads to
@@
We then suppose that the user chooses integers $n_{\ell}$ summing to $n=2^{m}$ that are nearly proportional to $\alpha_{\ell}^{2/(\rho+1)}$. The conservative criterion 1 leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho+2)}$.
For both criteria higher convergence rates $\rho$ tend to equalize the $n_{\ell}$. We can think of $n_{\ell}=n/L$ as arising in the limit as $\rho\to\infty$. We will also see equal $n_{\ell}$ in a minimax argument in Section 5. The conservative criterion 1 favors stratum sizes more nearly equal than under criterion 0 while not exactly equal as the even more conservative minimax argument gives.
We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for criteria 0 and $1$, and three different hypotheses on $\rho$. They lead to the four different rates given in Table 1. The case for criterion 0 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation. The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$. This is also true for the plain MC rate ($\rho=1$) under the conservative criterion 1.
When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then our criteria do not distinguish between different values of $\rho$. After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
When we do have information about the relative sizes of $\tau_{\ell}$, then we can replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}$ in the formulas in Table 1. In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934). When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$, then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}c_{\ell}^{-1/2}$ in Table 1. When using this extra information, we should order the strata to be in non-increasing order of $\alpha_{\ell}\tau_{\ell}^{1/2}c_{\ell}^{-1/2}$ for either criterion.
@@
Table 1: This table shows ideal rates for $n_{\ell}$ as functions of $\alpha_{\ell}$ with $C_{0}$ and $C_{1}$ for criteria 0 and 11 respectively. These rates assume the variance within stratum $\ell$ is proportional to $n_{\ell}^{-\rho}$.
[…]
@@ 3.2 Suboptimality
Under criterion 0, we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
@@
The factor $n^{-\rho}$ has canceled from the numerator and denominator and so we get the same rate of convergence with an incorrect $\gamma$ as we would with the optimal $\gamma=\rho$. Only the lead constant is affected. The analogous formula for criterion 1 is
$I_{1}(\gamma\!\mid\!\rho)$ $=\frac{\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}^{1-\rho/(\gamma+2)}\bigr)^{2}\bigl(\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr)^{\rho}}{\bigl(\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr)^{\rho+2}}.$
Under criterion 0 a cautious choice is to take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma_{0}+1)}$ where
@@
Next we consider recommendations that can be made generally without specific values of $\alpha_{\ell}$ in mind. The first thing to notice is that both $I_{0}$ and $I_{1}$ contain a numerator sum that could potentially have a negative power of $\alpha_{\ell}$. For general use where some $\alpha_{\ell}$ might be small, we want to avoid that possibility. There is no possibility of an even more negative power in the denominator to compensate. To keep a positive exponent in the numerator of $I_{0}$ we must have $2-2\rho/(\gamma+1)\geqslant 0$. After rearrangement this becomes a constraint that $\gamma\geqslant\rho-1$, whatever the unknown $\rho$ might be. When $\rho\doteq 3$ is a possibility, then we should take $\gamma\geqslant 2$. If instead we use $\gamma=1$ in a $\rho=3$ setting then the first factor in $I_{0}$ is greater than $\alpha_{L}^{-1}$. For $I_{1}$ we want $1-\rho/(\gamma+2)\geqslant 0$ or $\gamma\geqslant\rho-2$. Since we already had $\gamma\geqslant\rho-1$ using criterion 0, this is not a further constraint. When we know that $\rho\leqslant 2$, then any $\gamma\geqslant 1$ avoids a negative exponent for $\alpha_{\ell}$.
@@ 4 Making each $n_{\ell}$ a power of two
Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to a given total sample size $n=2^{m}$. We assume throughout that $n\geqslant L$ and that we want every $n_{\ell}\geqslant 1$. The advantage of having every $n_{\ell}$ be a power of two is that we then get scrambled net sampling within each of the $L$ strata as shown by Proposition 2. Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen 2020). The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
@@
The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc. 2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al. (2024)). The latter reference mentions that the number of solutions is asymptotic to $R\theta^{L}$ for $R\doteq 0.14$ and $\theta\doteq 1.794$. This is a somewhat faster growth rate than the Fibonacci numbers have: their counterpart to $\theta$ is $(1+\sqrt{5})/2\doteq 1.618$. For $L=10$, the approximation is close to 48.3.
@@
[…]
Under criterion 0, the ideal proportion of samples in stratum $\ell$ is
@@
which we consider to be the most ‘under-represented’ of the eligible strata. The analogous algorithm under criterion $1$ is to double the sample size of stratum
@@
The pseudo-code for forward allocation algorithm for criterion 0 is given in Algorithm 1.
@@ 5 Minimax choice
$\Delta_{N}$ $=\biggl\{\bm{n}=(n_{1},\dots,n_{L})\in\mathbb{N}^{L}\Bigm|\sum_{\ell=1}^{L}n_{\ell}=N\biggr\},\quad\text{and}\quad$
$\mathcal{T}$ $=\Bigl\{\bm{\tau}=(\tau_{1},\dots,\tau_{L})\in[0,\infty)^{L}\bigm|\max_{1\leqslant\ell\leqslant L}\tau_{\ell}>0\Bigr\}$
@@
We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used. For criteria 0 and $1$ we define suboptimality ratios
@@
[…]